At the end of the current year, the accounts receivable account has a debit balance of...

  1. At the end of the current year, the accounts receivable account has a debit balance of $1,206,000 and sales for the year total $13,680,000.

    1. The allowance account before adjustment has a credit balance of $16,300. Bad debt expense is estimated at 1/4 of 1% of sales.
    2. The allowance account before adjustment has a credit balance of $16,300. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $52,200.
    3. The allowance account before adjustment has a debit balance of $9,200. Bad debt expense is estimated at 1/2 of 1% of sales.
    4. The allowance account before adjustment has a debit balance of $9,200. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $76,400.

    Determine the amount of the adjusting entry to provide for doubtful accounts under each of the assumptions (a through d) listed above.

    a. $
    b. $
    c. $
    d. $
Answer #1

Determine the amount of the adjusting entry to provide for doubtful accounts under each of the assumptions (a through d) listed above.

a. 13680000*1%*1/4 = 34200
b. 52200-16300 = 35900
c. 13680000*1%*1/2 = 68400
d. 76400+9200 = 85600